OpenStudy (anonymous):

find the slope of the line using the two plotted points in the graph. (8,-2) (-6,5)

OpenStudy (nenadmatematika):

-1/2

OpenStudy (anonymous):

How did you get _1/2

OpenStudy (anonymous):

-1/2

OpenStudy (nenadmatematika):

(5-(-2))/(-6-8)=7/(-14)=-1/2

OpenStudy (anonymous):

Thank you for showing me